Consider the following boundary value problem, x?y + 13xy + (36+2) y = 0, y(1) = 0, yler/8 ) = 0 (a) Find the eigenvalues.
Answer #1

э + 13x. + (36 +d) a чCi) - 11 8 ح yle о Let z = ln x dy he Х dx dz Dy x² d²y ау dzz de (Т-Ф)) = (у – dz so УСр - 1) 9 + 3 +1) 9 – 9 – - 67. ЕК се t - 6z Cate — 9 - Схь + C₂ X Inx 1) с. + C₂ mx 9 yli)= с . о (? с – о 9(ews ) с. (73) о 6 Сухо (ев )Now 2) ma -6 Į ki y y= ë bz [ a cos(lez) + Q Sin (42)] ē y=x6 [ acos (Klmx) + C Sim (klmx)] y= сі yle) = 1 18 (6) [ acos (17)